# The Orientation Boundary for Step-Duplicating Recursors: Mechanized Impossibility, Escape, and Certification

### 19 Mar 2026 v7

@@ 5.3 Root-step confluence via Newman
The side condition $\kappa^{M}(t)=\varnothing$ on R_merge_cancel is part of the guarded confluence design, not a cosmetic strengthening: when $\kappa^{M}(t)\neq\varnothing$, the collapsing root step is blocked, and the proof switches to the blocked-guard case localJoin_merge_cancel_guard_kappa_ne. The collapsing case itself is handled separately by localJoin_merge_tt. This coupling between the termination measure component $\kappa^{M}$ and the confluence guard is a deliberate design choice: the guarded fragment uses the same multiset infrastructure for both termination (the second component of $\mu^{3}_{c}$) and confluence (the merge-cancel guard). In KO7 the same multiset is exactly what distinguishes harmless collapsing cases from forks that would reintroduce the bad payload, so separating the two would duplicate the same structural case split rather than simplify it.
@@ 5.4 Ordinal upper-bound calibration
This yields a mechanized order-type isomorphism: $(\mathrm{Multiset}\;\mathbb{N},\mathrm{DM})\cong\omega^{\omega}$ (cf. [8]). The file now packages the exact statement and the cofinal bridge under the theorem names phaseB_cnf_scaffold_exact_order_type and phaseB_cnf_scaffold_cofinal.
@@
Proposition 5.13 (Derivation-length bound for SafeStepCtx).
For any term $t$ of structural size $n$, any SafeStepCtx reduction chain from $t$ has length at most $\mathrm{towerBound}(n)$, where $\mathrm{towerBound}$ is the explicit tower-exponential function
$\mathrm{towerBound}(0)=6,\qquad\mathrm{towerBound}(k{+}1)=2^{2\cdot\mathrm{towerBound}(k)+5}+\mathrm{towerBound}(k)+1.$
Machine-checked in Meta/SafeStep_Complexity_Ordinal.lean as safestep_length_bounded_by_size. The proof factors through the computable numeric witness ctxFuel, which strictly decreases on every context-closed safe step. This bound is numerically much cruder than the ordinal calibration at $\omega^{\omega}\!\cdot\!2$: it gives a concrete tower-style majorant rather than a sharp fast-growing or slow-growing extraction. The technical gap is that ctxFuel absorbs context descent by explicit exponentiation, whereas a genuine Moser–Weiermann extraction would need a direct fundamental-sequence control argument for the calibrated ordinal measure itself. A tighter Moser–Weiermann-style descent argument remains future work.
@@
Machine-checked in Meta/SafeStep_Complexity_FastGrowing.lean as fastGrow, fgOmegaEnvelope, towerBound_le_fgOmegaEnvelope, and safestep_length_bounded_by_fgOmegaEnvelope. This does not yet constitute a full Moser–Weiermann slow-growing extraction from the ordinal embedding, but it replaces the old tower-only presentation with an explicit finite-level fast-growing-family envelope aligned with the $\omega^{\omega}$ calibration.
@@ 6 External validation: TTT2 and CeTA
We submitted the full KO7 system (8 rules, no SafeStep guards) to TTT2 [17]. TTT2’s MAYBE results are tool search failures, not proofs of impossibility; the YES result below independently corroborates the Lean theorem 5.4. Here the TPDB/TRS semantics is the usual context-closed rewriting semantics, so the CeTA certificate corresponds to Lean’s StepCtxFull, not merely the root relation Step.
@@
The three YES results use modular or path-order methods; the five MAYBE results use direct/global methods (polynomial, KBO, matrix, forward/backward instantiation). The barrier theorems (Section 4) cover eight scalar numeric direct classes, additive, transparent-compositional, affine, restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized bounded polynomial, and max-plus, together with one symbolic variable-condition barrier, an explicit KBO corollary, and four vector / pair extensions, tracked fixed-dimension componentwise, tracked lexicographic dimension-$2$, balanced mixed-coordinate dimension-$2$ componentwise, and weighted scalar-projection componentwise. The KBO MAYBE outcome is explained directly by Corollary 4.15. The POLY MAYBE outcome is theorem-covered for the formalized direct polynomial families: affine, restricted quadratic, bounded cross-term, bounded multilinear, generalized bounded polynomial, and max-plus are all blocked, while the successful nonlinear witness $W$ lies outside those direct classes by cross-variable coupling and non-transparency. The MAT(d)(d) MAYBE outcomes now split cleanly. Fixed-row tracked-component and row-sum weighted-projection interpretations are theorem-covered by the existing matrix barriers and the explicit corollaries in Meta/MatrixProjectionCoverage.lean; only arbitrary mixed matrix interpretations remain outside the present theorem stack. The FBI MAYBE entry remains outside the theorem layer here: no general forward/backward-instantiation impossibility result is claimed. CPF certificates and CeTA logs are in Artifacts/ttt2/. The Lean module Meta/TPDB_Export.lean renders the exact KO7 TTT2 input and proves that the generated TPDB text matches the artifact text embedded in that module, which in turn corresponds to Artifacts/ttt2/KO7_full_step.trs. The artifact also includes a runnable verifier, lakeexeverifyTpdbExport, that checks the embedded text against the on-disk .trs file. In addition, Meta/TTT2_CertificateReplay.lean replays the mathematical core of the FAST certificate inside Lean: rather than parsing the external CPF, it re-establishes the relevant mathematical content under Lean definitions, namely the single recursive SCC, the projected argument (tool index $2$, paper index $3$), and the resulting well-founded DP pair proof.
@@
All three YES strategies received a certified verdict from CeTA 2.36. The four MAYBE strategies were correctly rejected. These external tool runs corroborate the method separation predicted by the barrier theorems. They are not, however, part of the formal impossibility proof: the MAYBE outcomes are search failures of bounded heuristic strategies, not theoretical impossibility results.
@@ 7 Formalization structure
  • Operational boundary tools: computable barrier witnesses, including extended extractors for quadratic, max-plus, and projected matrix-side subclasses, a small synthesis-oracle wrapper around them, and a decidable coefficient-table classifier for the formalized scalar and tracked-primary pair families.
@@
  • Complexity and ordinal layers: guarded-context ctxFuel termination, contextual derivation-length bounds, exact merge-spine normalizer cost, and ordinal upper/lower calibration.
  • Catalog, SCC, and ablation layers: theorem-backed depth and pure-precedence families, verified excluded-method results, the delayed-duplication and multiplicity-preserving SCC theorems, the linear-recursion ablation, and the object-level collapse surrogate.
  • Artifact bridge: Lean-side TPDB export for the full KO7 root TRS, an exact-match theorem for the checked TTT2 artifact text, a narrow internal replay of the FAST DP/subterm certificate core, and a runnable verifier for the on-disk .trs file.
@@
SafeStepCtx wf_SafeStepCtxRev, confluentSafeCtx, safestep_length_bounded_by_size, safestep_length_bounded_by_fgOmegaEnvelope Internal Lean proof Partial context closure of the guarded system
@@
The formalization uses three proof shapes: (i) direct full-step orientation (polynomial $W$, KO7-specialized MPO), stated on the unrestricted Step relation; (ii) guarded-fragment certification (triple-lex measure, confluence, normalizer), on SafeStep; and (iii) transformed-relation arguments (dependency pairs), factored through a small reusable fragment for projection-rank descent and SCC-path reasoning. The reusable schema barrier layer is now also exposed through a dedicated public API module, OperatorKO7/SchemaAPI.lean, so downstream developments can import the generic barrier stack and witness/classifier tooling without pulling in the full KO7 certification chain. The support layer is also computational: Meta/BarrierClass_Classifier.lean classifies coefficient tables into the formalized barrier families, Meta/BarrierWitness.lean and Meta/BarrierWitness_Extended.lean return certified violating triples across the scalar and projected matrix-side subclasses, and Meta/SynthesisOracle.lean packages them into a small oracle interface for direct-orientation attempts. All declarations are sorry-free. The termination, normalization, and barrier developments are computationally effective; the ordinal-calibration layer (Meta/DM_OrderType.lean, Meta/DM_OrderType_LowerBound.lean) uses standard classical infrastructure from Lean 4 / Mathlib (Classical.choice, propext) for noncomputable ordinal reasoning. For reviewability, the repository now ships a bidirectional proof-dependency graph in Docs/KO7_BLUEPRINT.md and Docs/ko7_blueprint.json, mapping paper theorem labels to primary Lean declarations and back.
@@
From a proof perspective, the development is not a Lean transcription of the paper statements. Across more than sixty Lean/build files, its main theorem layer consists of impossibility theorems, computed barrier witnesses, boundary classifiers, and a small internal replay of the FAST certificate core, all inside one Lean library, rather than reconstruction of externally generated positive certificates. Mathlib supplies the well-foundedness background for Dershowitz–Manna multisets and the ambient ordinal infrastructure used in Meta/DM_OrderType.lean, but the precise $\omega^{\omega}$ image and lower-bound argument in Meta/DM_OrderType_LowerBound.lean and the full context-closed lift of the nonlinear witness through Meta/ContextClosed_SN_Full.lean are built in this repository. Those two modules are the clearest mechanization bottlenecks: the former because it must recover a sharp ordinal calibration rather than only well-foundedness, and the latter because the polynomial witness has to be threaded through every constructor context together with an explicit derivation-length control.
@@
The paper does not prove arbitrary-target reachability decidability, an exact order type for the full triple-lex image of guarded reductions, or generic MPO/LPO metatheory over arbitrary signatures. It also does not claim that the computable triple-lex certificate handles unrestricted Step, nor that Meta/KBO_Impossible.lean formalizes generic KBO metatheory; the full root Step arguments are instead the nonlinear witness and KO7-specialized MPO proofs, while the KBO corollary is extracted from the symbolic variable-condition abstraction. Nor does it lift the tree-based barrier stack wholesale to higher-order rewriting; the current artifact instead isolates a sharing-aware surrogate in Meta/SharingBarrierLift.lean showing exactly where the no-sharing assumption is load-bearing. Each of those boundaries is explicit in the corresponding sections below.
@@
For reproducibility, the active artifact is pinned to Lean toolchain leanprover/lean4:v4.22.0-rc4 and Mathlib commit 632465e4b02cb70a5dfa4cfe15468e8a62c2bd85; lake-manifest.json freezes the transitive dependency graph. The repository-level note artifact/REPRODUCIBILITY.md documents the minimal replay path, and scripts/make_referee_bundle.py produces a referee-facing bundle containing the frozen Lean snapshot, generated documentation, and archived Artifacts/ttt2/ validation trail together with replay instructions. Separately, artifact/TPDB_SUBMISSION_PREP.md and scripts/stage_tpdb_submission.py stage, but do not perform, a future TPDB submission.

### 08 Apr 2026 v8

@@ 5.3 Root-step confluence via Newman
The side condition $\kappa^{M}(t)=\varnothing$ on R_merge_cancel is part of the guarded confluence design, not a cosmetic strengthening: when $\kappa^{M}(t)\neq\varnothing$, the collapsing root step is blocked, and the proof switches to the blocked-guard case localJoin_merge_cancel_guard_kappa_ne. The collapsing case itself is handled separately by localJoin_merge_tt. This coupling between the termination measure component $\kappa^{M}$ and the confluence guard is a deliberate design choice: the guarded fragment uses the same multiset infrastructure for both termination (the second component of $\mu^{3}_{c}$) and confluence (the merge-cancel guard). In KO7 the same multiset is exactly what distinguishes harmless collapsing cases from forks that would reintroduce the bad payload, so separating the two would duplicate the same structural case split rather than simplify it. Concretely, if $\mathrm{merge}\;t\;t\to t$ were allowed while $t$ still contained an active $\mathrm{rec}\Delta$ subterm, the two copies could unfold differently before any later merge, producing distinct descendants instead of a trivially joinable collapse.
@@ 5.4 Ordinal upper-bound calibration
This yields a mechanized order-type isomorphism: $(\mathrm{Multiset}\;\mathbb{N},\mathrm{DM})\cong\omega^{\omega}$ (cf. [9]). The exact statement and the cofinal bridge appear in the theorem names phaseB_cnf_scaffold_exact_order_type and phaseB_cnf_scaffold_cofinal.
Proposition 5.12 (Realized lower families for the full $\mu_{3c}$ image).
For every $n\in\mathbb{N}$, there exist traces $t_{0},t_{1}$ such that
$\omega^{n+2}\;\leq\;\texttt{lex3cToOrd}(\mu_{3c}(t_{0}))\;<\;\omega^{\omega}$
and
$\omega^{\omega}+\omega^{n+2}\;\leq\;\texttt{lex3cToOrd}(\mu_{3c}(t_{1}))\;<\;\omega^{\omega}\cdot 2.$
Formalized in Meta/Mu3c_Image_LowerBound.lean as payloadTower_mu3c_lower, payloadTower_mu3c_lt_block_cap, flaggedTower_mu3c_lower, flaggedTower_mu3c_lt_two_block_cap, mu3c_lower_family_block0, and mu3c_lower_family_block1. This does not upgrade the full $\mu_{3c}$ story to an exact order-type theorem, but it does show that the realized trace image is nontrivial in both $\delta$-blocks and reaches arbitrarily high ordinals below the established $\omega^{\omega}\!\cdot\!2$ cap.
@@
Proposition 5.14 (Single-exponential derivation-length bound for SafeStepCtx).
For any term $t$ of structural size $n$, any SafeStepCtx reduction chain $t\to^{m}u$ satisfies
$m+1\;\leq\;2^{2n}.$
[…]
Proposition 5.15 (Explicit exponential lower family for SafeStepCtx).
Let
$t_{0}:=\mathrm{merge}\;\mathrm{void}\;\mathrm{void},\qquad t_{n+1}:=\mathrm{rec}\Delta\;\mathrm{void}\;t_{n}\;(\delta(\delta\,\mathrm{void})).$
Then there exist terms $u_{n}$ and exact contextual lengths $m_{n}$ such that
$\texttt{SafeStepCtxPow}\;m_{n}\;t_{n}\;u_{n},\qquad 2^{n}\leq m_{n},$
$\mathrm{termSize}(t_{n})=5n+3,\qquad\mathrm{termSize}(u_{n})=2^{n+2}-3.$
Hence the guarded context-closed derivational complexity is genuinely single-exponential in structural size, up to constant factors in the exponent, and the same family exhibits explicit exponential normal-form size growth from linear-size guarded inputs.
The Lean proof is in Meta/SafeStepCtx_Complexity_LowerBound.lean; the relevant identifiers are ctxLowerFamily, ctxLowerNF, ctxLowerLen, termSize_ctxLowerFamily, termSize_ctxLowerNF, ctxLowerFamily_pow_nf, two_pow_le_ctxLowerLen, safeStepCtx_has_singleExponential_lower_family, safeStepCtx_has_singleExponential_size_family, and safeStepCtx_has_singleExponential_output_family. Together with Proposition 5.14, this pins the guarded contextual derivation-length story between matching single exponentials: the old ctxFuel/towerBound theorem remains available as a coarse explicit envelope, but the actual certified complexity class is now exponential rather than merely bounded above by a tower. The same family also provides a direct duplication-growth witness relevant to bounded-growth heuristics such as match-bound-style approaches, without claiming any generic match-bound metatheory.
@@
See Meta/SafeStep_Complexity_FastGrowing.lean (fastGrow, fgOmegaEnvelope, towerBound_le_fgOmegaEnvelope, and safestep_length_bounded_by_fgOmegaEnvelope). This remains a coarse explicit corollary to Proposition 5.14. Proposition 5.15 shows that the actual guarded-context complexity is already exponential, so the fast-growing envelope should be read only as a retained explicit majorant. The paper-facing guarded-context hierarchy theorems are the two Cichon-style propositions below.
Proposition 5.17 (Cichon-style upper bound for SafeStepCtx).
Define the contextual notation
$\mathrm{mwCtxNote}(t)\;=\;\mathrm{lex3Note}\bigl(\mathrm{deltaFlag}(t),(\kappa^{M}(t),\texttt{ctxFuel}(t))\bigr)$
and the associated hierarchy bound
$\mathrm{mwCtxBound}(t)\;=\;\mathrm{cichon}\bigl(\mathrm{mwCtxNote}(t),0\bigr).$
Then for every SafeStepCtx reduction chain $t\to^{n}u$,
$n\;\leq\;\mathrm{mwCtxBound}(t).$
Moreover $\mathrm{mwCtxNote}(t)<\omega^{\omega}\!\cdot\!2$ for every $t$.
Proved in Meta/SafeStep_Complexity_MW_Ctx.lean (mwCtxNote, mwCtxBound, mwCtxNote_lt_opow_omega_mul_two, ctxFuel_le_mwCtxBound, safestepctx_length_le_mwCtxBound, ctxOrdinalObstruction_rootNote_increases, and not_all_safeStepCtx_rootNote_drop). This is a conservative Moser–Weiermann lift for the context-closed relation: it packages the already-certified contextual potential ctxFuel into the finite tail of the same lex3Note family used by the root calibration, yielding a genuine cichon theorem while keeping the bound in the calibrated $\omega^{\omega}\!\cdot\!2$ block. The same file also proves that the exported root-side calibrated note can increase on a concrete SafeStepCtx step, so Proposition 5.17 should be read as a theorem-backed conservative lift rather than as a direct contextual fundamental-sequence descent theorem.
Proposition 5.18 (Direct exact contextual extraction on a separate note package).
Define the exact contextual note
$\mathrm{ctxExactNote}(t)\;=\;\mathrm{lex3Note}\bigl(0,(0,\texttt{ctxFuel}(t))\bigr)$
and the associated hierarchy bound
$\mathrm{ctxExactBound}(t)\;=\;\mathrm{cichon}\bigl(\mathrm{ctxExactNote}(t),0\bigr).$
Then for every SafeStepCtx reduction chain $t\to^{n}u$,
$n\;\leq\;\mathrm{ctxExactBound}(t).$
Moreover $\mathrm{ctxExactNote}(t)<\omega$ for every $t$, and every contextual chain induces an exact fundamental-sequence descent on this new note family with total length $\texttt{ctxFuel}(t)-\texttt{ctxFuel}(u)$.
The Lean development is in Meta/SafeStep_Complexity_MW_CtxExact.lean (ctxExactNote, ctxExactBound, ctxExactNote_lt_omega, safeStepCtx_exact_step, safeStepCtxPow_exact_drop, and safestepctx_length_le_ctxExactBound). This gives the direct contextual exact-control theorem on a separate contextual note/control package. It does not reuse the root-calibrated note; that impossibility is recorded in Proposition 5.17. Instead, it turns the already-certified contextual potential into a standalone exact fundamental-sequence package.
@@ 6 External validation: TTT2 and CeTA
We submitted the full KO7 system (8 rules, no SafeStep guards) to TTT2 [18]. TTT2’s MAYBE results are tool search failures, not proofs of impossibility; the three YES results below independently corroborate the Lean theorem 5.4. Here the TPDB/TRS semantics is the usual context-closed rewriting semantics, so the CeTA certificate corresponds to Lean’s StepCtxFull, not merely the root relation Step.
@@
The three YES results use modular or path-order methods; the five MAYBE results use direct/global methods (polynomial, KBO, matrix, forward/backward instantiation). The barrier theorems (Section 4) cover eight scalar numeric direct classes, additive, transparent-compositional, affine, restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized bounded polynomial, and max-plus, together with one symbolic variable-condition barrier, an explicit KBO corollary, and four vector / pair extensions, tracked fixed-dimension componentwise, tracked lexicographic dimension-$2$, balanced mixed-coordinate dimension-$2$ componentwise, and weighted scalar-projection componentwise. The KBO MAYBE outcome is explained directly by Corollary 4.16. The POLY MAYBE outcome is theorem-covered for the formalized direct polynomial families: affine, restricted quadratic, bounded cross-term, bounded multilinear, generalized bounded polynomial, and max-plus are all blocked, while the successful nonlinear witness $W$ lies outside those direct classes by cross-variable coupling and non-transparency. The MAT(d)(d) MAYBE outcomes now split cleanly. Fixed-row tracked-component and row-sum weighted-projection interpretations are theorem-covered by the existing matrix barriers and the explicit corollaries in Meta/MatrixProjectionCoverage.lean; only arbitrary mixed matrix interpretations remain outside the present theorem stack. The FBI MAYBE entry remains outside the theorem layer here: no general forward/backward-instantiation impossibility result is claimed. CPF certificates and CeTA logs are in Artifacts/ttt2/. The Lean module Meta/TPDB_Export.lean renders the exact KO7 TTT2 input and proves that the generated TPDB text matches the artifact text embedded in that module, which in turn corresponds to Artifacts/ttt2/KO7_full_step.trs. The artifact also includes a runnable verifier, lakeexeverifyTpdbExport, that checks the embedded text against the on-disk Artifacts/ttt2/KO7_full_step.trs file. In addition, Meta/TTT2_CertificateReplay.lean is a narrow Lean-side replay of the mathematical core of the FAST certificate: rather than parsing the external CPF, it re-establishes under Lean definitions the single recursive SCC, the projected argument (tool index $2$, paper index $3$), and the resulting well-founded DP pair proof.
@@
All three YES strategies received a certified verdict from CeTA 2.36. The four MAYBE strategies were rejected in the narrow certification sense that no checkable termination certificate was produced for CeTA to replay; this is not a proof of impossibility. These external tool runs corroborate the method separation predicted by the barrier theorems. They are not, however, part of the formal impossibility proof: the MAYBE outcomes are search failures of bounded heuristic strategies, not theoretical impossibility results.
@@ 7 Formalization structure
  • Operational boundary tools: computable barrier witnesses, including extended extractors for quadratic, max-plus, and projected matrix-side subclasses, theorem-backed canonical witness budgets for those extractors, a small synthesis-oracle wrapper around them, and a decidable coefficient-table classifier for the formalized scalar and tracked-primary pair families.
@@
  • Complexity and ordinal layers: guarded-context ctxFuel termination, a new single-exponential contextual derivation-length theorem with matching explicit lower family, exact merge-spine normalizer cost, and ordinal upper/lower calibration.
  • Catalog, SCC, and ablation layers: theorem-backed depth and pure-precedence families, verified excluded-method results, the delayed-duplication SCC theorem together with its finite cyclic, raw-graph, automatic closed-path / round-trip extraction, and bounded finite-round-trip reachability extensions, the multiplicity-preserving SCC theorem together with its finite cyclic, raw-graph, automatic closed-path / round-trip extraction, and bounded finite-round-trip reachability extensions, the linear-recursion ablation, and the object-level collapse surrogate.
  • Artifact bridge: Lean-side TPDB export for the full KO7 root TRS, an exact-match theorem for the checked TTT2 artifact text, a narrow internal replay of the FAST DP/subterm certificate core, and a runnable verifier for the on-disk Artifacts/ttt2/KO7_full_step.trs file.
Table 10: Formalization layers, primary declarations, and evidence types in the active artifact.
@@
SafeStepCtx wf_SafeStepCtxRev, confluentSafeCtx, safeStepCtx_length_le_contextualExpBound, safestep_length_bounded_by_fgOmegaEnvelope Internal Lean proof Partial context closure of the guarded system; single-exponential structural bound with matching explicit lower family plus coarser inherited envelopes
@@
The formalization uses three proof shapes: (i) direct full-step orientation (polynomial $W$, KO7-specialized MPO), stated on the unrestricted Step relation; (ii) guarded-fragment certification (triple-lex measure, confluence, normalizer), on SafeStep; and (iii) transformed-relation arguments (dependency pairs), factored through a small reusable fragment for projection-rank descent and SCC-path reasoning. The reusable schema barrier layer is also exposed through a dedicated public API module, OperatorKO7/SchemaAPI.lean, so downstream developments can import the generic barrier stack and witness/classifier tooling without pulling in the full KO7 certification chain.
The support layer is computational as well. Meta/BarrierClass_Classifier.lean classifies coefficient tables into the formalized barrier families. Meta/BarrierWitness.lean and Meta/BarrierWitness_Extended.lean return certified violating triples across the scalar and projected matrix-side subclasses. Meta/BarrierWitness_Budgets.lean packages the corresponding canonical construction budgets, and Meta/SynthesisOracle.lean wraps the witness layer in a small oracle interface for direct-orientation attempts. All declarations are sorry-free.
The termination, normalization, and barrier developments are computationally effective. The ordinal and hierarchy-calibration layer (Meta/DM_OrderType.lean, Meta/DM_OrderType_LowerBound.lean, Meta/SafeStep_Complexity_MW_Root.lean, Meta/SafeStep_Complexity_MW_Ctx.lean, and Meta/MPO_ProofTheoreticBound.lean) uses standard classical infrastructure from Lean 4 / Mathlib (Classical.choice, propext) for noncomputable ordinal reasoning. For reviewability, the repository ships a bidirectional proof-dependency graph in Docs/KO7_BLUEPRINT.md and Docs/ko7_blueprint.json, mapping paper theorem labels to primary Lean declarations and back.
@@
From a proof perspective, the development is not a Lean transcription of the paper statements. Across more than sixty Lean/build files, the main theorem layer consists of impossibility theorems, computed barrier witnesses, boundary classifiers, and a small internal replay of the FAST certificate core, all inside one Lean library rather than as reconstructions of externally generated positive certificates. Mathlib supplies the well-foundedness background for Dershowitz–Manna multisets and the ambient ordinal infrastructure used in Meta/DM_OrderType.lean, but the precise $\omega^{\omega}$ image and lower-bound argument in Meta/DM_OrderType_LowerBound.lean and the full context-closed lift of the nonlinear witness through Meta/ContextClosed_SN_Full.lean are built in this repository.
The repository also includes an upstream-prep staging surface for the general DM package in Meta/DM_UpstreamSurface.lean. This is a preparation artifact for a future Mathlib contribution, not a claimed upstream merge. Those two modules are the clearest mechanization bottlenecks: the former because it must recover a sharp ordinal calibration rather than only well-foundedness, and the latter because the polynomial witness has to be threaded through every constructor context together with an explicit derivation-length control.
@@
The paper does not prove arbitrary-target reachability decidability, an exact order type for the full triple-lex image of guarded reductions, or generic MPO/LPO metatheory over arbitrary signatures. It also does not claim that the computable triple-lex certificate handles unrestricted Step, nor that Meta/KBO_Impossible.lean formalizes generic KBO metatheory; the full root Step arguments are instead the nonlinear witness and KO7-specialized MPO proofs, while the KBO corollary is extracted from the symbolic variable-condition abstraction. Likewise, Meta/WPO_PolynomialBarrier.lean is not generic WPO metatheory: it only lifts the bounded direct-polynomial barrier to direct orders whose strict comparison is certified by that same algebraic branch. Nor does it lift the tree-based barrier stack wholesale to higher-order rewriting; the current artifact instead isolates a sharing-aware surrogate in Meta/SharingBarrierLift.lean showing exactly where the no-sharing assumption is load-bearing. Each of those boundaries is explicit in the corresponding sections below.
For reproducibility, the active artifact is pinned to the following Lean / Mathlib environment:
Lean toolchain: leanprover/lean4:v4.22.0-rc4
Mathlib commit: 632465e4b02cb70a5dfa4cfe15468e8a62c2bd85
@@
The file lake-manifest.json freezes the transitive dependency graph. Public artifact-facing notes and staged outputs are collected under the single repository root Artifacts/. The repository-level note Artifacts/REPRODUCIBILITY.md documents the minimal replay path. Artifacts/MICRO_BENCHMARKS.md records representative local build/check times together with the archived TTT2 run times. The script scripts/make_referee_bundle.py writes a referee-facing bundle in the public Artifacts/ tree, containing the frozen Lean snapshot, generated documentation, and the archived Artifacts/ttt2/ validation trail together with replay instructions. Separately, scripts/stage_tpdb_submission.py stages, but does not perform, a future TPDB submission in that same Artifacts/ tree.
Trusted-base note. Beyond ordinary elaboration, the intended replay path is lake build OperatorKO7, lake exe verifyTpdbExport, and #print axioms on headline declarations after import OperatorKO7. In the pinned environment, the following spot-check commands return only the standard Lean / Mathlib axioms propext, Classical.choice, and Quot.sound:
OperatorKO7.MetaCM.wf_SafeStepRev_c
OperatorKO7.PolyInterpretation.wf_StepRev_poly
OperatorKO7.MetaDM.dm_order_type_omega_omega
MetaSN_KO7.safeStepCtx_length_le_contextualExpBound
The active tree currently uses native_decide 19 times: 15 concrete finite sanity checks in the dependency-pair extraction infrastructure, 3 small controlled hierarchy checks in Meta/OrdinalHierarchy_Controlled.lean, and 1 literal TPDB-text equality check in Meta/TPDB_Export.lean. No generic CPF parser, external solver oracle, or user-declared axiom layer is part of the main theorem surface; the only noncomputable proof layer is the explicitly identified classical ordinal calibration infrastructure above.